OpenStudy (cliffordtapp):

if points (1,4) and (4,2) line on line k. what is the slope of the line that perpendicular to k.

OpenStudy (anonymous):

First you have to find the slop of the slop of k. You find this from: \[\frac{ \Delta y }{ \Delta x }\] As slope (or gradient) is just how much y changes with respect to x. Therefore the slop is =(4-2)/(1-4) =2/-3 =-2/3 If a line is perpendicular to another the gradients of both lines multiply to -1 (you can look up the reason why if you want its pretty simple) -2/3 * (gradient of perpindicularline) = -1 gradient of perpindicualr line = 3/2

OpenStudy (anonymous):

wow is said slop twice :P i mean slope

OpenStudy (anonymous):

And i just mean slope of k, not slope of slope of k. My grammar failed hard